Hello my problem is:
The average arithmetic mean of two positive numbers, x and y, is 10. If the smaller number is subtracted from twice the larger number, the result is 14. If y>x, which of the following pairs of equations could correctly express the information above?
So I know the total numbers are 2, which is x and y. And the average is 10. Now I'm stuck as to do what to do next.

Answer by Alan3354(69443)

The average arithmetic mean of two positive numbers, x and y, is 10. If the smaller number is subtracted from twice the larger number, the result is 14. If y>x, which of the following pairs of equations could correctly express the information above?
So I know the total numbers are 2, which is x and y. And the average is 10. Now I'm stuck as to do what to do next.
=========
If the avg is 10, and there are 2 numbers, the sum is 20
---
y + x = 20
2y- x = 14
------------- Add
3y = 34
y = 34/3
x = 26/3